# VibES: Induced Vibration for Persistent Event-based Sensing

### v1

@@ Abstract
Event cameras are a bio-inspired class of sensors that asynchronously measure per-pixel intensity changes. Under fixed illumination conditions in static or low-motion scenes, rigidly mounted event cameras are unable to generate any events, becoming unsuitable for most computer vision tasks. To address this limitation, recent work has investigated motion-induced event stimulation that often requires complex hardware or additional optical components. In contrast, we introduce a lightweight approach to sustain persistent event generation by employing a simple rotating unbalanced mass to induce periodic vibrational motion. This is combined with a motion-compensation pipeline that removes the injected motion and yields clean, motion-corrected events for downstream perception tasks. We demonstrate our approach with a hardware prototype and evaluate it on real-world captured datasets. Our method reliably recovers motion parameters and improves both image reconstruction and edge detection over event-based sensing without motion induction.
@@ 1 Introduction
Figure 1: Qualitative illustration of our method. No Vibration. With a static event camera and no induced motion, the accumulated event image appears blurred and lacks sharp edges, while the yytt slice shows little temporal structure. With Vibration. Introducing controlled vibrations stimulates the sensor, increasing the number of events and producing sinusoidal traces in the yytt slice. We develop an Extended Kalman Filter (EKF) to track the vibrational motion in a region of the scene (red). With Vibration + Compensation. We use the EKF-tracked region to estimate and compensate for the sinusoidal motion across the entire scene. The motion-compensated accumulated image recovers sharp structures, and the yytt slice aligns with the stable motion-compensated track (magenta), revealing the underlying scene. An inset (bottom left) shows the reference scene displayed in front of the camera with induced motion.
Most computer vision algorithms rely on conventional image sensors — such as RGB or grayscale cameras — that capture synchronous image frames at fixed intervals. While widely adopted, these frame-based sensors suffer from well-known limitations, including motion blur, latency due to frame timing, and high power consumption [19]. These drawbacks become particularly problematic in robotics applications where fast dynamics, low latency, and energy efficiency are critical, such as in autonomous navigation [41], feature tracking [32], and real-time scene understanding [27].
Event cameras have emerged as a promising alternative. These bio-inspired sensors asynchronously detect per-pixel changes in log-intensity, producing a sparse stream of events with microsecond latency and high dynamic range [19]. Event cameras inherently avoid motion blur, operate at lower power [20], and enable low-latency perception, making them well suited for resource-constrained, high-speed robotics. Recent advances in event-based feature tracking and visual odometry have demonstrated strong performance in dynamic scenarios [32, 45], further highlighting their potential for robust autonomy.
However, a fundamental limitation persists: event generation depends on motion. In static scenes — or when edges align with the direction of the camera motion — events are not triggered. This results in event sparsity, loss of spatial information, and perceptual fading [7], which undermines long-term feature tracking, edge detection, and high-quality image reconstruction [24]. Consequently, the performance of event-based systems degrades significantly without sufficient camera or scene motion.
@@ 2 Related Work
Event cameras are increasingly explored in robotics and computer vision for their ability to capture scene dynamics with high temporal precision. In this section, we review their role in perception, outline complementary hybrid sensing strategies, and discuss microsaccade-inspired methods for maintaining event generation in otherwise static scenes.
@@
Event cameras have gained significant attention in robotics and computer vision communities due to their unique capabilities, including high dynamic range, low latency, and motion-blur-free imaging. Recent methods have exploited these properties for feature tracking in challenging conditions [21, 32, 51].
Common applications of event cameras in robotics include visual localization [46], edge detection [5], and motion compensation [18]. Due to their high temporal resolution, frequency estimation with event cameras has also been studied extensively [37, 3, 42]. Other notable applications include image reconstruction [45, 26, 23] and image deblur [36, 25]. We show that for motion compensation, edge detection and image reconstruction tasks, our method can provide a high-quality information stream, taking advantage from our induced motion.
@@
Beyond microsaccades, other strategies for event generation include moving cameras at high speeds for 3D reconstruction [29], pairing event cameras with conventional frames for static localization [11], and employing visual tracking to follow objects [22]. Machine learning-based approaches have also been developed to infer object motion from sparse event data [6], with some relying on continuous-motion assumptions, such as in human choreography capture applications [49]. Our system does not rely on conventional image frames or on the assumption of fast motion in the scene. Instead, we propose a fully model-based event-centric approach that exploits the intrinsic properties of events and their sensitivity to motion, generating a continuous stream of information. This stream can be leveraged for reconstruction, edge detection, and the tasks discussed above, even in otherwise static scenes.
@@
Inspired by human fixational eye movements (FEMs), researchers have explored creating both saccadic and non-saccadic artificial FEMs to address the limitations of event cameras in static scenes. Proposed mechanisms include pan–tilt tables [50, 43, 44, 34, 13], mirrors [30], a rotating wedge-prism system [24], and a rotating quarter-wave plate in front of the camera [31]. However, these approaches require precisely constructed mechanical systems combined with per-system calibration. Our mechanical system is comparatively simple, and our processing pipeline can automatically adjust to different system parameters. This makes our approach more appealing for real-world use, where systems that require careful tuning are less favourable than those that adjust parameters online. Concurrent work has attempted to recover motion from unstructured camera jitter emulating spacecraft vibration [2], using a clustering-based algorithm. However, the absence of a prior on the camera’s jitter reduces reconstruction accuracy, leading to errors of several pixels in image space.
Most similar to our approach is that of Botao et al. [24], which proposes a rotating wedge prism to generate events continuously and utilizes a position-encoding hardware to track and remove the induced motion, thereby stabilizing texture and enhancing information output. Our method introduces a mechanically simpler approach, requiring no custom optical elements or position encoders such as in [24]. Instead, we directly estimate the state of the sinusoidal motion from the event stream itself and perform parameter estimation online. Moreover, we demonstrate that our software stack is agnostic to the underlying hardware. On data generated by the rotating wedge-prism in [24], we show that our system performs equally well in terms of motion compensation quality and event stream consistency. Finally, our experiments show that VIBES enables the same downstream use cases proposed in  [24] — particularly in image reconstruction and edges extraction — with simpler hardware and a more general motion-tracking framework.
@@ 3 Methodology
Our approach aims to stimulate the pixels of an event camera to produce a continuous stream of events, enabling high-quality image reconstruction and sharp edge detection from raw event data. We first describe the mechanical system responsible for generating vibrational motion and its effect on event generation in the camera (Sec. 3.1). We then present the motion compensation pipeline that allows us to recover the underlying scene structure (Sec. 3.2). Finally, we describe the physical prototype implementation of our system (Sec. 3.3).
@@ 3.1 Camera Motion Model
Figure 2: Schematic of the mass–spring–damper model. Camera setup at two time steps, $t_{0}$ a) and $t_{1}$ b). The rotation of an off-axis mass $m$ with angular velocity $\omega$ induces planar displacements $\Delta x,\Delta y$ (only the vertical component is illustrated). The estimated oscillation frequency, $\frac{2\pi}{T}$, corresponds to the motion perceived by the camera. Note that this differs from the natural frequency of the off-axis mass $\omega$, due to inertial effects.
@@
where $\hat{A}=\sqrt{\hat{A}_{X}^{2}+\hat{A}_{Y}^{2}}$ is the oscillation amplitude, with $\hat{A}_{X}$ and $\hat{A}_{Y}$ as the components along $\mathbf{x}$- and $\mathbf{y}$-axis in camera coordinates. As such, the motion of the camera when viewing a static scene can be expressed sinusoidally. Further details are provided in Supplementary Sec. S1.1.
The camera motion is described by introducing a virtual camera frame, denoted as $\mathcal{C}$, shown in Fig. 3. In our formulation, $\mathcal{C}$ remains static and acts as the reference frame, while the physical camera, $\mathcal{C}^{\prime}$, moves along a controlled, circular trajectory in the xxyy image plane of $\mathcal{C}$.
@@
Thus from Eq. 5, we identify three key unknowns for motion compensation in one axis: the amplitude $A_{y}$ (dependent on scene depth and oscillation magnitude), the phase $\phi_{Y}$, and the coordinate $v$ in the virtual frame. Accurate estimation of these parameters allows precise inversion of the apparent motion in the event stream, recovering the underlying scene as if the camera were stationary.
@@ 3.2 Motion Compensation Setup
Once a model of the induced motion has been established, the next step is to eliminate motion from the incoming event stream. This requires estimating the dynamic motion parameters, which varies as the camera or the scene changes over time. To achieve this, we employ a series of software modules, illustrated in Fig. 4. Below, we describe each stage of the pipeline and its role in accurate motion compensation.
@@ 3.2.2 Frequency Estimation with NUFFT
Since the camera motion is oscillatory, recovering its frequency is essential. However, events are sampled irregularly in time, which makes standard Fourier analysis unsuitable. To obtain an initial estimate of the angular velocity parameter in Eq. 5, we employ the non-uniform fast Fourier Transform (NUFFT) [4]. The NUFFT decomposes a signal into its constituent frequencies but is specifically designed to operate on irregularly sampled data — making it ideal for event-based inputs, which are inherently asynchronous and non-uniform in time. Before applying the NUFFT, we remove the DC component from the tracker signal (Sec. 3.2.1) to isolate oscillatory motion from the static offset introduced by the patch location. Temporal samples $t_{i}$ are then rescaled to $\pi,\pi]$, yielding normalized timestamps $\bar{t}_{i}$. This produces a zero-mean, time-normalized samples ${(\bar{u^{\prime}}_{i},\bar{v^{\prime}}_{i},\bar{t}_{i})}_{i=1}^{N}$ which are passed to the NUFFT. The transform returns the top $M$ dominant frequencies, from which we extract the angular frequency $\hat{\omega}$ of the camera motion. Estimation is performed independently for the $X$ and $Y$ axes. Using $\hat{\omega}$, we fit the tracked samples to a sinusoidal model using least-squares optimization, recovering the amplitudes and phase offsets of the motion. These parameters initialize the Extended Kalman Filter (EKF) described in Sec. 3.2.3.
@@ 3.2.3 Extended Kalman Filter Setup
The Extended Kalman Filter (EKF) refines the motion parameters over time by incorporating a dynamical model and accounting for observation noise. This yields temporally consistent parameter estimates, which are essential for reliable motion compensation. We model all motion parameters dynamically in the EKF, but they are updated differently. Oscillatory frequency and phase are assumed to be globally constant across the image plane, whereas the oscillation amplitude depends on the scene depth. We make the assumption that all objects in the same scene share the same depth plane, although our implementation allows for multiple trackers which may have regions with different depths. Thus, we vary the $u^{\prime},v^{\prime}$ parameters with each specific tracker to allow for tracking at different depths.
@@
$v^{\prime}=a_{y}\sin(\theta)+b_{y}\cos(\theta)+v$ (7)
@@
This state is updated as new tracker measurements arrive, with the corresponding Jacobians provided in Supplementary Appendix S2. The filter can operate on the entire image plane or, more effectively, on localized patches centered around feature points. Our implementation allows multiple trackers to be instantiated in parallel, enabling flexible motion parameter estimation across the scene.
@@ 3.3 Hardware Prototype
For all real-world data collection and experiments, we use the Prophesee EVK3 event camera with IMX636 (1280x720 px) sensor. To induce motion in the Prophesee camera, we rigidly attach a DC motor [14] to the body of the event camera. Affixed to the DC motor shaft is a mass, which has a slight eccentricity. This mass was custom machined, although any off-axis weight with sufficient eccentricity and mass would suffice. A casing enclosing the event camera was 3D printed, and wrapped in foam. The foam acts as a passive-spring damper, enabling the motion model described in Sec. 3.1. Lastly, a 3D printed shell supporting the foam and event camera casing was used to ensure 2D planar motion for the event camera and to enable mounting for real-world use. We refer to Supplementary Sec. S2.3 for more details regarding our prototype and modeling assumptions.
@@ 4.1 Data Collection
We record real-world scenes and produce output data for three settings. First, we capture the scene using a standard event camera (S-EV). Second, we capture the scene with vibration (V-EV). Finally, we apply our motion compensation algorithm to the acquired V-EV data (VIBES). Overall, we evaluated our method on four real-world scenes: AMI-EV [24], Logo, Pattern Checkerboard, and Pattern. Each scene consists of a textured pattern moving in front of the camera. The Logo pattern contains text and rich textures, moving slowly while performing partial rotations. The Pattern Checkerboard and Pattern sequences involve planar patterns moving at different speeds, with the Pattern exhibiting faster motions. Images of all patterns are provided in Supplementary Sec. S2.4. To ensure a fair comparison in the real-world setting, we employed a Franka robotic arm to repeatably move various patterns in front of the camera for the V-EV and S-EV settings. We mounted the camera on the physical prototype mass-spring-damper system described in Sec. 3.3.
@@ 4.2 Metrics
We evaluated the performance of our method for image quality, edges, and texture detection using the following metrics. We provide mathematical definitions of all metrics in Supplementary Sec. S2.5.
@@
  • Natural Image Quality Evaluator (NIQE) [33]. We compute NIQE on reconstructed images obtained using E2VID [38, 39] with a 10 ms event stream window. Lower values indicate reconstructions closer to the statistical properties of natural images and higher perceptual quality.

### v2

@@ Abstract
Event cameras are a bio-inspired class of sensors that asynchronously measure per-pixel intensity changes. Under fixed illumination conditions in static or low-motion scenes, rigidly mounted event cameras are unable to generate any events and become unsuitable for most computer vision tasks. To address this limitation, recent work has investigated motion-induced event stimulation, which often requires complex hardware or additional optical components. In contrast, we introduce a lightweight approach to sustain persistent event generation by employing a simple rotating unbalanced mass to induce periodic vibrational motion. This is combined with a motion-compensation pipeline that removes the injected motion and yields clean, motion-corrected events for downstream perception tasks. We develop a hardware prototype to demonstrate our approach and evaluate it on real-world datasets. Our method reliably recovers motion parameters and improves both image reconstruction and edge detection compared to event-based sensing without motion induction.
Supplementary Material: For code and data please visit https://papers.starslab.ca/vibes/.
@@ 1 Introduction
Figure 1: Qualitative illustration of our method. (a) No Vibration. With a static event camera and no induced motion, the accumulated event image appears blurred and lacks sharp edges, while the yytt slice shows little temporal structure. (b) With Vibration. Introducing controlled vibrations stimulates the sensor, increasing the number of events and producing sinusoidal traces in the yytt slice. We develop an Extended Kalman Filter (EKF) to track the vibrational motion in a region of the scene (red). (c) With Vibration + Compensation. We use the EKF-tracked region to estimate and compensate for the sinusoidal motion across the entire scene. The motion-compensated accumulated image recovers sharp structures, and the yytt slice aligns with the stable motion-compensated track (magenta), revealing the underlying scene. An inset (bottom left) shows the reference scene displayed in front of the camera with induced motion.
Most computer vision algorithms rely on conventional image sensors—such as RGB or greyscale cameras—that capture synchronous image frames at fixed intervals. While widely adopted, these frame-based sensors suffer from well-known limitations, including motion blur, latency due to frame timing, and high power consumption [19]. These drawbacks are especially problematic in robotics applications where fast dynamics, low latency, and energy efficiency are critical, for example in navigation [41], feature tracking [32], and real-time scene understanding [27].
Event cameras have emerged as a promising alternative. These bio-inspired sensors asynchronously detect per-pixel changes in log-intensity, producing a sparse stream of events with microsecond latency and high dynamic range [19]. Event cameras inherently reduce motion blur, operate at lower power [20], and enable low-latency perception, making them well-suited for resource-constrained, high-speed robots.
However, event cameras do have a fundamental limitation: event generation depends on motion. In static scenes, or when edges align with the direction of the camera motion, events are not triggered. This results in event sparsity, loss of spatial information, and perceptual fading [7], which undermines long-term feature tracking, edge detection, and high-quality image reconstruction [24]. Consequently, the performance of event-based systems degrades significantly without sufficient camera or scene motion.
@@ 2 Related Work
Event cameras are increasingly used in robotics and computer vision due to their high temporal resolution and ability to capture fast scene dynamics. In this section, we review their role in perception, outline complementary hybrid sensing strategies, and discuss microsaccade-inspired methods for maintaining event generation in otherwise static scenes.
@@
Event cameras offer several advantages over conventional cameras, including high dynamic range and microsecond-level temporal resolution. These properties have been exploited for feature tracking under challenging conditions [21, 32, 50].
In robotics, event cameras have been applied to visual localization [45], edge detection [5], and motion compensation [18], frequency estimation [37, 3, 42], image reconstruction [45, 26, 23] and deblurring [36, 25]. We show that for motion compensation, edge detection, and image reconstruction tasks, our method can provide a high-quality information stream, taking advantage of the induced motion.
@@
Beyond microsaccades, other strategies for event generation include moving stereo event cameras at high speed for 3D reconstruction [29], pairing event cameras with conventional frames for static localization [11], and employing visual tracking for object following [22]. Machine learning-based approaches have also been developed to infer object motion from sparse event data [6], with some methods relying on continuous-motion assumptions, for example in human choreography capture applications [48]. Our system does not rely on conventional image frames or on the assumption of fast motions in the scene. Instead, we propose a fully model-based event-centric approach that exploits the intrinsic properties of events and their sensitivity to motion, generating a continuous stream of information. This stream can be leveraged for reconstruction, edge detection, and the tasks discussed above, even in otherwise static scenes.
@@
Inspired by fixational eye movements in human vision, several works have proposed artificial microsaccades to enable event generation in static scenes. Existing solutions rely on mechanically induced motion using pan–tilt units [49, 43, 44, 34, 13], mirrors [30], rotating wedge prisms [24], or polarization-based optical elements [31]. These systems typically require specialized hardware and careful calibration. Our approach employs a simpler mechanical setup and a processing pipeline that adapts automatically to system parameters, making it more suitable for real-world deployment. Concurrent work has explored motion recovery from unstructured camera jitter inspired by spacecraft vibration [2], but the lack of a motion prior limits reconstruction accuracy.
The approach most similar to our own is that of He et al. [24]. They employ a rotating wedge prism to generate events continuously and utilize position-encoding hardware to track and remove the induced motion. Our method introduces a mechanically simpler approach, requiring no custom optical elements or position encoders such as in [24]. Instead, we directly estimate the state of the sinusoidal motion from the event stream itself and perform parameter estimation online. Moreover, we demonstrate that our software stack is agnostic to the underlying hardware. On data generated by the rotating wedge-prism in [24], we show that our system performs equally well in terms of motion compensation quality and event stream consistency. Finally, our experiments show that our system enables the same downstream use cases proposed in  [24]—particularly in image reconstruction and edges extraction—with simpler hardware and a more general motion-tracking framework.
@@ 3 Methodology
Our approach aims to excite the pixels of an event camera to produce a continuous stream of events, enabling high-quality image reconstruction and sharp edge detection from raw event data. We first describe the mechanical system responsible for generating vibrational motion and its effect on event generation in the camera (Sec. 3.1). We then present the motion compensation pipeline that allows us to recover the underlying scene structure (Sec. 3.2). Finally, we describe our physical hardware prototype (Sec. 3.3).
@@ 3.1 Camera Motion Model
Figure 2: Schematic of the mass–spring–damper model. Camera setup at two time steps, (a) $t_{0}$ and (b) $t_{1}$. The rotation of an off-axis mass $m$ with angular velocity $\omega$ induces planar displacements $\Delta x,\Delta y$ (only the vertical component is illustrated). The estimated oscillation frequency, $\frac{2\pi}{T}$, corresponds to the motion perceived by the camera. Note that this differs from the natural frequency of the off-axis mass $\omega$, due to inertial effects.
@@
where $\hat{A}=\smash{\sqrt{\hat{A}_{X}^{2}+\hat{A}_{Y}^{2}}}$ is the oscillation amplitude, with $\hat{A}_{X}$ and $\hat{A}_{Y}$ as the components along $\mathbf{x}$- and $\mathbf{y}$-axis in camera coordinates. As such, the motion of the camera when viewing a static scene follows a sinusoid. Further details are provided in the supplementary material, Sec. S1.1.
The camera motion is described by introducing a virtual camera frame, denoted as $\mathcal{C}$, shown in Fig. 3. In our formulation, $\mathcal{C}$ remains static and acts as the reference, while the physical camera, $\mathcal{C}^{\prime}$, moves along a controlled, circular trajectory in the xxyy image plane of $\mathcal{C}$.
@@
From Eq. 5, we identify three key unknowns for motion compensation along one axis: the amplitude $A_{y}$ (dependent on scene depth and oscillation magnitude), the phase $\phi_{Y}$, and the coordinate $v$ in the virtual frame. Accurate estimation of these parameters allows precise inversion of the apparent motion in the event stream and recovery of the underlying scene as if captured by a stationary camera.
@@ 3.2 Motion Compensation Setup
Once a model of the induced motion has been established, the next step is to eliminate motion from the incoming event stream. This requires estimating the dynamic motion parameters, which vary as the camera or the scene changes over time. To achieve this, we employ a series of software modules, illustrated in Fig. 4. Below, we describe each stage of the pipeline and its role in accurate motion compensation.
@@ 3.2.2 Frequency Estimation with NUFFT
Since the camera motion is oscillatory, recovering its frequency is essential. However, events are sampled irregularly in time, which makes standard Fourier analysis unsuitable. To obtain an initial estimate of the angular velocity parameter in Eq. 5, we employ the non-uniform fast Fourier transform (NUFFT) [4]. The NUFFT decomposes a signal into its constituent frequencies but is specifically designed to operate on irregularly sampled data—making it ideal for event-based inputs that are inherently asynchronous and non-uniform in time. Before applying the NUFFT, we remove the DC component from the tracker signal (Sec. 3.2.1) to isolate oscillatory motion from the static offset introduced by the patch location. Temporal samples $t_{i}$ are then rescaled to $\pi,\pi]$, yielding normalized timestamps $\bar{t}_{i}$. This produces a zero-mean, time-normalized samples ${(\bar{u^{\prime}}_{i},\bar{v^{\prime}}_{i},\bar{t}_{i})}_{i=1}^{N}$ which are passed to the NUFFT. The transform returns the top $M$ dominant frequencies, from which we extract the angular frequency $\hat{\omega}$ of the camera motion. Estimation is performed independently for the $X$ and $Y$ axes. Using $\hat{\omega}$, we fit the tracked samples to a sinusoidal model using least-squares optimization, recovering the amplitudes and phase offsets of the motion. These parameters initialize the extended Kalman filter (EKF) described in Sec. 3.2.3.
@@ 3.2.3 Extended Kalman Filter Setup
An extended Kalman filter (EKF) refines the motion parameters over time by incorporating a dynamical model and accounting for observation noise. This yields temporally consistent parameter estimates, which are essential for reliable motion compensation. We model all motion parameters dynamically in the EKF, but they are updated differently. Oscillatory frequency and phase are assumed to be globally constant across the image plane, whereas the oscillation amplitude depends on the scene depth. We make the assumption that all objects in the same scene share the same depth plane, although our implementation allows for multiple trackers which may have regions with different depths. Thus, we vary the $u^{\prime},v^{\prime}$ parameters with each specific tracker to allow for tracking at different depths.
@@
$v^{\prime}=a_{y}\sin(\theta)+b_{y}\cos(\theta)+v,$ (7)
@@
This state is updated as new tracker measurements arrive, with the corresponding Jacobians provided in Supplementary Sec. S2. The filter can operate on the entire image plane or, more effectively, on localized patches centered around feature points. Our implementation allows multiple trackers to be instantiated in parallel, enabling flexible motion parameter estimation across the scene.
@@ 3.3 Hardware Prototype
For all real-world data collection and experiments, we use a Prophesee EVK3 event camera equipped with an IMX636 sensor (1280 × 720 pixels). To induce motion, we rigidly attach a DC motor [14] to the camera body and mount an off-centre mass on the motor shaft. The mass is custom-machined; however, any off-axis weight with sufficient eccentricity and mass is sufficient to generate the desired motion.
The event camera is enclosed in a 3D-printed casing wrapped in foam, which acts as a passive spring–damper system and enables the motion model described in Sec. 3.1. The case constrains the motion to a 2D plane and provides a rigid interface for real-world mounting. Additional details on the mechanical prototype are provided in the supplementary material (Sec. S2.3).
@@ 4.1 Data Collection
We recorded real-world scenes and produced output data for three settings. First, we captured the scene using a standard event camera (S-EV). Second, we capture the scene with vibration (V-EV). Finally, we apply our motion compensation algorithm to the acquired V-EV data (VibES). Overall, we evaluated our method on four real-world scenes: AMI-EV [24], Logo, Pattern Checkerboard, and Pattern. Each scene consists of a textured pattern moving in front of the camera. The Logo pattern contains text and rich textures, moving slowly while performing partial rotations. The Pattern Checkerboard and Pattern sequences involve planar patterns moving at different speeds, with the Pattern exhibiting faster motions. Images of all patterns are provided in Supplementary Sec. S2.4. To ensure a fair comparison in the real-world setting, we employed a Franka robotic arm to move various patterns in front of the camera repeatably for the V-EV and S-EV settings. We mounted the camera on the physical prototype mass-spring-damper system described in Sec. 3.3.
@@ 4.2 Metrics
We evaluated the performance of our method for image quality, edges, and texture detection using the following metrics. We provide mathematical definitions of all metrics in our supplementary material, Sec. S2.5.
@@
  • Natural Image Quality Evaluator (NIQE) [33]. We compute NIQE on reconstructed images from E2VID [38, 39] with a 10 ms event stream window. Lower values indicate reconstructions closer to the statistical properties of natural images and higher perceptual quality.